Statement related to the next three questions:

You are single and have an adjusted gross income of $35,000 and taxable income of $24,000. Let’s assume that, according to the marginal tax rate schedule for single filers, you would owe 10% on your first $8,925 of taxable income, plus 15% of the remaining taxable income.

8) What is your marginal tax rate? A) 0% B) 9% C) 13.1% D) 15% E) 20.1%

9) What is the average tax rate you paid on your taxable income? A) 0% B) 9% C) 13.1% D) 15% E) 20.1%

10) What is the average tax rate you paid on your adjusted gross income? A) 0% B) 9% C) 13.1% D) 15% E) 20.1%
Answer #1

Answer 8:

Correct answer is:

D) 15%

Explanation:

Given:

According to the marginal tax rate schedule for single filers, you would owe 10% on your first $8,925 of taxable income, plus 15% of the remaining taxable income.

As such marginal tax rate is 15%.

Hence option D is correct and other options A, B, C and E are incorrect

Answer 9:

Correct answer is:

C) 13.1%

Explanation:

Given:

According to the marginal tax rate schedule for single filers, you would owe 10% on your first $8,925 of taxable income, plus 15% of the remaining taxable income.

Taxable income = $24,000

Hence:

Tax = 8925 * 10% + 15% * (24000 - 8925) = 3153.75

Average tax = Tax / Taxable income = 3153.75 /24000 =13.1%

Hence option C is correct and other options A, B, D and E are incorrect.

Answer 10:

Correct answer is:

B) 9%

Explanation:

Given:

Adjusted gross income = $35,000

Average tax rate you paid on your adjusted gross income = 3153.75 / 35000 = 9%

Hence option B is correct and other options A, C, D and E are incorrect.